A rectangular room is 10 m wide and 6m high. If the area of the 4 walls is 540m2, its length will be
a) 40m c) 45m b) 70m d) 35m
1
Expert's answer
2013-01-10T09:21:46-0500
S = high*2*wide + high*2*length
S - area of the 4 walls
high - high of the room
wide - wide of the room
length - length of the room
length = S/(2*high) -wide = 540/12 - 10 = 45 -10 =35
Answer: d) 35 m